# An efficient probabilistic hardware architecture for diffusion-like models

### v1

@@ Abstract
The proliferation of probabilistic AI has promoted proposals for specialized stochastic computers. Despite promising efficiency gains, these proposals have failed to gain traction because they rely on fundamentally limited modeling techniques and exotic, unscalable hardware. In this work, we address these shortcomings by proposing an all-transistor probabilistic computer that implements powerful denoising models at the hardware level. A system-level analysis indicates that devices based on our architecture could achieve performance parity with GPUs on a simple image benchmark using approximately 10,000 times less energy.
@@
The unprecedented recent investment in large-scale AI systems will soon put a strain on the world’s energy infrastructure. Every year, U.S. firms spend an amount larger than the inflation-adjusted cost of the Apollo program on AI-focused data centers [17, 65]. By 2030, these data centers could consume 10% of all of the energy produced in the U.S. [4].
@@
Despite this enormous bet on scaling today’s AI systems, they may be far from optimal in terms of energy efficiency. Existing AI systems based on autoregressive large language models (LLMs) are valuable tools in white-collar fields [45, 41, 50, 51, 12, 52], and are being adopted by consumers faster than the internet [8]. However, LLMs were architected specifically for GPUs [68], hardware originally intended for graphics, whose suitability for machine learning was discovered accidentally decades later [18, 15].
Figure 1: Leveraging CMOS probabilistic hardware in ultra-efficient AI systems. The central result of this article: an all‑transistor probabilistic computer running a denoising thermodynamic model (DTM) could match GPU performance on a simple modeling benchmark while using about $\numprint{10000}\times$ less energy. All models are trained on binarized Fashion-MNIST [72] and evaluated with Fréchet Inception Distance (FID) [32]. DTM variants are of increasing depth, chaining 2–8 sequential Energy-Based Models (EBMs). GPU baselines cover single‑step VAE [42] and GAN [31], plus DDPM [62] at varying numbers of steps. We also compare DTM to a monolithic EBM across multiple mixing‑time limits. The horizontal axis shows the energy needed for generating a single new image using the trained model (inference).
Had a different style of hardware been popular in the last few decades, AI algorithms would have evolved in a completely different direction, and possibly a more energy-efficient one. This interplay between algorithm research and hardware availability is known as the "Hardware Lottery" [37], and it entrenches hardware-algorithm pairings that may be far from optimal.
@@
Therefore, prudent planning calls for systematic exploration of other types of AI systems in search of energy-efficient alternatives. Active efforts include mixed‑signal compute‑in‑memory accelerators [5], photonic neural networks [7], and neuromorphic processors that emulate biological spiking [30, 59].
The development of more efficient computers for AI is challenging because it requires innovation not only at the component level but also at the system level. It is insufficient to invent a new technology that performs a mathematical operation efficiently in isolation; one must also know how to combine multiple components to run a practical algorithm. In addition to these integration challenges, GPU performance per joule is doubling every few years [66], making it very difficult for cutting-edge computing schemes to gain mainstream adoption.
@@
Probabilistic computing is an attractive approach because it can connect directly to AI at the system level via Energy-Based Models (EBMs). EBMs are a well-established model class in contemporary deep learning and have been competitive with the state of the art in tasks like image generation and robotic path planning [63, 39].
@@
Hardware implementations of EBMs work with special model families that adhere to physical constraints such as locality, sparsity, and connection density. Thanks to these constraints, probabilistic computers can utilize specialized stochastic circuitry to efficiently and quickly produce samples from a Boltzmann distribution [61]. Depending on the precise kind of hardware being used, this sampling may occur as part of the natural dynamics of the device [53, 70, 3, 67, 25] or may be orchestrated using an algorithm like Gibbs sampling [49, 10, 60, 57]. Using probabilistic hardware to accelerate EBMs falls under the broad umbrella of thermodynamic computing [19].
@@
In this work, we address these issues and propose a commercially viable probabilistic computing system. Our contributions extend from broad architectural choices down to designing and fabricating novel mixed-signal RNG circuitry.
At the top level, we introduce a new probabilistic computer architecture that runs Denoising Thermodynamic Models (DTMs) instead of monolithic EBMs. As their name suggests, rather than using the hardware’s EBM to model data distributions directly, DTMs sequentially compose many hardware EBMs to model a process that denoises the data gradually. Diffusion models [62, 35] also follow this denoising procedure and are much more capable than EBMs. This key architectural change addresses a fundamental issue with previous approaches and represents the first scalable method for applying probabilistic hardware to machine learning.
Additionally, we show that our new architecture can be implemented at scale using present-day CMOS processes by experimentally demonstrating an all-transistor RNG that is fast, energy efficient, and small. By using transistors as the only building blocks of our RNG circuits, we avoid the significant and ambiguous communication overheads that can occur at interfaces between technologies. Furthermore, the absence of such communication overhead allows for the principled forecasting of device performance that we present in this work. Our RNG leverages the stochastic dynamics of subthreshold transistor networks, which we have recently studied in detail in Ref. [26].
Our system-level analysis indicates that combining our new architecture with our all-transistor probabilistic computing circuitry could achieve unprecedented energy efficiency in probabilistic modeling. Figure 1 compares the predicted performance and energy consumption of such a system to several standard deep-learning algorithms running on GPUs and a traditional EBM-based probabilistic computer. The DTM-based probabilistic computer system achieves performance parity with the most efficient GPU-based algorithm while using around four orders of magnitude less energy.
The remainder of this article will substantiate the results presented in Fig. 1, which are based on a combination of measurements from real circuits, physical models, and simulations. To begin, we introduce a fundamental compromise inherent in using EBMs as standalone models of data, which we refer to as the mixing-expressivity tradeoff. We then discuss how this compromise can be avoided by wielding EBMs as part of a denoising process rather than monolithically. Next, we outline how to build a hardware system using DTMs to implement this denoising process at a very low level. Then, we study simulations of this hardware system, further justifying the results shown in Fig. 1 and highlighting some of the practical merits of DTMs compared to existing approaches. Finally, we conclude by discussing how the capabilities of probabilistic accelerators for machine learning may be scaled by merging them with traditional neural networks.
@@ I The Challenge with EBMs
Fitting an MEBM corresponds to assigning low energies to values of $x$ where data are abundant and high energies to values of $x$ that are far from data. Real-world data are often clustered into distinct modes [20, 9], meaning that a MEBM that fits data well will have a complex, rugged energy landscape with many deep valleys surrounded by tall mountains. This complexity is illustrated by the cartoon in Fig. 2 (a).
@@
Specifically, sampling algorithms that operate in high dimensions (such as Gibbs sampling [48]) are locally-informed iterative procedures, meaning that they sample a landscape by randomly making small movements in the space based on low-dimensional information. When using such a procedure to sample from Eq. (1), the probability that the iteration will move up in energy to some state $X[k+1]$ is exponentially small in the energy increase compared to the current state $X[k]$, i.e.,
@@
For large differences in energy, like those encountered when trying to move between two valleys separated by a significant barrier, this probability can be very close to zero. These barriers grind the iterative sampler to a halt.
The mixing-expressivity tradeoff (MET) summarizes this issue with existing probabilistic computer architectures, reflecting the fact that modeling performance and sampling hardness are coupled for MEBMs. Specifically, as the expressivity (modeling performance) of an MEBM increases, its mixing time (the amount of computational effort needed to draw independent samples from the MEBM’s distribution) becomes progressively longer, resulting in expensive inference and unstable training [13, 21].
@@
Figure 2: The mixing-expressivity tradeoff. (a) A cartoon illustrating the mixing-expressivity tradeoff in EBMs. It shows a projection of an energy landscape fit to a simple dataset. The "airplane" mode is well separated from the "dog" mode, with very little data in between. Progressively better fits of the EBM to the data tend to feature larger energy barriers $\Delta E$ between the modes, making the EBM increasingly difficult to sample from. (b) An example of the effect of the mixing-expressivity tradeoff on model performance as measured using the Fashion-MNIST dataset. The blue curve in the plot shows the results of experiments on MEBMs with limited allowed mixing time. Performance and mixing time are strongly correlated. Mixing times were computed by fitting an exponential function to the large-lag behavior of the autocorrelation function; see Appendix K. In contrast, a DTM (orange cross) has higher performance despite substantially lower sampling requirements.
@@ II Denoising thermodynamic models
DTMs merge EBMs with diffusion models, offering an alternative path for probabilistic computing that assuages the MET. DTMs are a slight generalization of recent work from deep learning practitioners that has pushed the frontier of EBM performance [27, 74, 73, 75].
@@
In traditional diffusion models, the forward process is made to be sufficiently fine-grained (using a large number of steps $T$) such that the conditional distribution of each step in the reverse process takes some simple form (such as Gaussian or categorical). This simple distribution is parameterized by a neural network, which is then trained to minimize the Kullback-Leibler (KL) divergence between the joint distributions $Q$ and $P_{\theta}$,
@@
See Appendix A.A.2 for more details.
@@
Equation (7) allows for a compromise between the number of steps in the approximation to the reverse process and the difficulty of sampling at each step. As the number of steps in the forward process is increased, the effect of each noising step becomes smaller, meaning that $\mathcal{E}^{f}_{t-1}$ more tightly binds $x^{t}$ to $x^{t-1}$. This binding can simplify the distribution given in Eq. (7) by imposing an energy penalty that prevents it from being strongly multimodal; see Appendix A.A.4 for further discussion.
@@
Introducing latent variables allows the size and complexity of the probabilistic model to be increased independently of the data dimension.
@@
See Appendix A.A.6 for further details. Note that this property relies on the normalizing constant associated with the distribution in Eq. (6) being independent of $x^{t-1}$.
@@ III Denoising Thermodynamic Computers
Practical implementations of the DTCA utilize natural-to-implement EBMs that exhibit sparse and local connectivity, as is typical in the literature [49]. This constraint allows sampling of the EBM to be performed by massively parallel arrays of primitive circuitry that implement Gibbs sampling. Refer to Appendices B and C for a further theoretical discussion of the hardware architecture.
A key feature of the DTCA is that $\mathcal{E}^{f}_{t-1}$ can be implemented efficiently using our constrained EBMs. Specifically, for both continuous and discrete diffusion, $\mathcal{E}^{f}_{t-1}$ can be implemented using a single pairwise interaction between corresponding variables in $x^{t}$ and $x^{t-1}$; see Appendix A.A.1 and C.C.1 for details. This structure can be reflected in how the chip is laid out to implement these interactions without violating locality constraints.
Critically, Eq. (8) places no constraints on the form of $\mathcal{E}_{t-1}^{\theta}$. Therefore, we are free to use EBMs that our hardware implements especially efficiently. At the lowest level, this corresponds to high-dimensional, regularly structured latent variable EBM. If more powerful models are desired, these hardware latent-variable EBMs can be arbitrarily scaled by combining them into software-defined graphical models.
@@
Figure 4: A programmable source of random bits. (a) A laboratory measurement of the operating characteristic of our RNG. The probability of the output voltage signal being in the high state ($x=1$) can be programmed by varying an input voltage. The relationship between $\mathbb{P}(x=1)$ and the input voltage is well-approximated by a sigmoid function. The inset shows the output voltage signal as a function of time for different input voltages. (b) The autocorrelation function of the RNG at the unbiased point ($\mathbb{P}(x=1)=0.5$). The decay is approximately exponential with the rate $\tau_{0}\approx 100\text{ns}$. (c) Estimating the effect of manufacturing variation on RNG performance. Each point in the plot represents the results of a simulation of an RNG circuit with transistor parameters sampled according to a procedure defined by the manufacturer’s PDK. Each color represents a different process corner, each for which $\sim 200$ realizations of the RNG were simulated. The "typical" corner represents a balanced case, whereas the other two are asymmetric corners where the two types of transistors (NMOS and PMOS) are skewed in opposite directions. The slow NMOS and fast PMOS case is worst performing for us due to an asymmetry in our design.
@@
Implementing our proposed hardware architecture using Boltzmann machines is particularly simple. A device will consist of a regular grid of Bernoulli sampling circuits, where each sampling circuit implements the Gibbs sampling update for a single variable $x_{i}$. The bias of the sampling circuits (probability that it produces 1 as opposed to 1-1) is constrained to be a sigmoidal function of an input voltage, allowing the conditional update given in Eq. (11) to be implemented using a simple circuit that adds currents such as a resistor network (See Appendix D.D.1).
@@
Specifically, the EBMs employed in this work were sparse, deep Boltzmann machines comprising $L\times L$ grids of binary variables, where $L=70$ was used in most cases. Each variable was connected to several (in most cases, 12) of its neighbors following a simple pattern. At random, some of the variables were selected to represent the data $x_{t-1}$, and the rest were assigned to the latent variables $z_{t-1}$. Then, an extra node was connected to each data node to implement the coupling to $x_{t}$. See Appendix C for further details on the Boltzmann machine architecture.
Due to our chosen connectivity patterns, our Boltzmann machines are bipartite (two-colorable). Since each color block can be sampled in parallel, a single iteration of Gibbs sampling corresponds to sampling the first color block conditioned on the second and then vice versa. Starting from some random initialization, this block sampling procedure could then be repeated for $K$ iterations (where $K$ is longer than the mixing time of the sampler, typically $K\approx 1000$) to draw samples from Eq. (7) for each step in the approximation to the reverse process.
@@
To enable a near-term, large-scale realization of the DTCA, we leveraged the shot-noise dynamics of subthreshold transistors [26] to build an RNG that is fast, energy-efficient, and small. Our all-transistor RNG is programmable and has the desired sigmoidal response to a control voltage, as shown by experimental measurements in Fig. 4 (a). The stochastic voltage signal output from the RNG has an approximately exponential autocorrelation function that decays in around $100$ ns, as illustrated in Fig. 4 (b). As shown in Ref. [26], this time constraint is much larger than the lower limit imposed by the correlation time of the noise in our transistors. The RNG could, therefore, be made much faster via an improved design. Appendix J provides further details about our RNG.
@@
A practical advantage to our all-transistor RNG is that detailed and proven foundry-provided models can be used to study the effect of manufacturing variations on our circuit design. In Fig. 4 (c), we use this process development kit (PDK) to study the speed and energy consumption of our RNG as a function of both systematic inter-wafer skews to the transistor parameters (process corners) and the expected variation within a single chip. We find that the RNG works reliably despite these non-idealities, meaning it can readily be scaled to the massive grids required by the DTCA.
@@
where $E_{\text{rng}}$ comes from the data in Fig. 4 (c). The term $E_{\text{bias}}$ is estimated using a physical model of a possible biasing circuit, and $E_{\text{clock}}$ and $E_{\text{comm}}$ are derived from physical reasoning about the costs of the clock and inter-cell communications respectively. $K_{\text{mix}}$ is the number of sampling iterations required to satisfactorily mix the chain for inference, which is generally less than the number of iterations used during training. $K_{\text{mix}}=250$ was used for the DTM (See Appendix D.D.4), while the mixing time measured in Fig. 2 was used for the MEBM.
This model is approximate, but it captures the underlying physics of a real device and provides a reasonable order-of-magnitude estimate of the actual energy consumption. Generally, given the same transistor process we used for our RNG and some reasonable selections for other free parameters of the model, we can estimate $E_{\text{cell}}\approx 2\;\text{fJ}$. See Appendix D for an exhaustive derivation of this model.
We use a simple model for the energy consumption of the GPU that underestimates the actual values. We compute the total number of floating-point operations (FLOPs) required to generate a sample from the trained model and divide that by the FLOP/joule specification given by the manufacturer. See Appendix E for further discussion.
Figure 5: Detailed results on the Fashion-MNIST dataset. (a) Images generated by a denoising model. Here, to achieve better-looking images, several binary variables were combined to represent a single grayscale pixel. The noisiness of the grayscale levels is an artifact of our embedding method; see Appendix G. (b) An experiment showing how DTMs are more stable to train than MEBMs. Complementing DTMs with the ACP completely stabilizes training. For the DTMs, the maximum $r_{yy}[K]$ value over all the layers is shown. (c) The effect of scaling EBM complexity on DTM performance. The grid size $L$ was modified to change the number of latent variables compared to the (fixed) number of data variables. Generally, EBM layers with more connectivity and longer allowed mixing times can utilize more latent variables and, therefore, achieve higher performance.

### v2

@@ Abstract
The proliferation of probabilistic AI has prompted proposals for specialized stochastic computers. Despite promising efficiency gains, these proposals have failed to gain traction because they rely on fundamentally limited modeling techniques and exotic, unscalable hardware. In this work, we address these shortcomings by proposing an all-transistor probabilistic computer that implements powerful denoising models at the hardware level. A system-level analysis indicates that devices based on our architecture could achieve performance parity with GPUs on a simple image benchmark using approximately 10,000 times less energy.
@@
The unprecedented recent investment in large-scale AI systems will soon put a strain on the world’s energy infrastructure. Every year, U.S. firms spend an amount larger than the inflation-adjusted cost of the Apollo program on AI-focused data centers [16, 63]. By 2030, these data centers could consume around 10% of all of the energy produced in the U.S. [4]. At the same time, model sizes, dataset sizes, and deployment footprints continue to grow rapidly.
@@
Existing AI systems based on autoregressive large language models (LLMs) are valuable tools in white-collar fields [44, 40, 48, 49, 11, 50], and are being adopted by consumers faster than the internet [7]. However, LLMs were architected specifically for GPUs [66], hardware originally intended for graphics, whose suitability for machine learning was discovered accidentally decades later [17, 14]. Had a different style of hardware been popular in the last few decades, AI algorithms would likely have evolved in a different direction, and possibly a more energy-efficient one. This interplay between algorithm research and hardware availability is known as the “hardware lottery” [35], and it entrenches hardware–algorithm pairings that may be far from optimal.
@@
Therefore, prudent planning calls for systematic exploration of other types of AI systems in search of energy-efficient alternatives. Active efforts include mixed-signal compute-in-memory accelerators [5], photonic neural networks [6], and neuromorphic processors that emulate biological spiking [29, 57]. These approaches explicitly confront the fact that computation is a physical process and that respecting physical constraints can lead to qualitatively different designs.
@@
Probabilistic computing is an attractive approach because it can connect directly to AI at the system level via Energy-Based Models (EBMs). EBMs are a well-established model class in contemporary deep learning and have been competitive with the state of the art in tasks like image generation and robotic path planning [61, 37]. Using probabilistic hardware to accelerate EBMs falls under the broad umbrella of thermodynamic computing [18], in which computation is expressed in terms of stochastic dynamics on energy landscapes.
Figure 1: Leveraging CMOS probabilistic hardware in ultra-efficient AI systems. The central result of this article: an all-transistor probabilistic computer running a denoising thermodynamic model (DTM) could match GPU performance on a simple modeling benchmark while using about $\numprint{10000}\times$ less energy. All models are trained on binarized Fashion-MNIST [69] and evaluated with Fréchet Inception Distance (FID) [31]. DTM variants are of increasing depth, chaining 2–8 sequential Energy-Based Models (EBMs). GPU baselines cover single-step VAE [41] and GAN [30], plus DDPM [60] at varying numbers of steps. We also compare DTM to a monolithic EBM across multiple mixing-time limits. The horizontal axis shows the energy needed for generating a single new image using the trained model (inference).
@@
Hardware implementations of EBMs work with special model families that adhere to physical constraints such as locality, sparsity, and limited connection density. Thanks to these constraints, probabilistic computers can utilize specialized stochastic circuitry to efficiently and quickly produce samples from a Boltzmann distribution [59]. Depending on the precise kind of hardware being used, this sampling may occur as part of the natural dynamics of the device [51, 67, 3, 64, 24] or may be orchestrated using an algorithm like Gibbs sampling [47, 9, 58, 55].
@@
In this work, we address these issues and propose a commercially viable probabilistic computing system that is designed for thermodynamic hardware from the outset. At the modeling level, we introduce Denoising Thermodynamic Models (DTMs), which repurpose hardware EBMs as denoising steps rather than monolithic models of the data distribution, thereby mitigating the mixing–expressivity tradeoff that plagues monolithic EBMs. At the architectural level, we propose the Denoising Thermodynamic Computer Architecture (DTCA), which tightly integrates DTMs into probabilistic hardware built from sparse, locally connected Boltzmann machines and an all-transistor RNG. Finally, we perform a system-level energy analysis, using measurements from an experimental thermodynamic RNG chip together with physical models and simulations, and show that a device based on this architecture could match the performance of GPU-based generative models on binarized Fashion-MNIST (as measured by Fréchet Inception Distance) while using approximately 10,000 times less energy per generated sample.
The remainder of this article will substantiate the results presented in Fig. 1, which are based on a combination of measurements from real circuits, physical models, and simulations. To begin, we introduce a fundamental compromise inherent in using EBMs as standalone models of data, which we refer to as the mixing–expressivity tradeoff. We then discuss how this compromise can be avoided by wielding EBMs as part of a denoising process rather than monolithically. Next, we outline how to build a hardware system using DTMs to implement this denoising process at a very low level. Then, we study simulations of this hardware system, further justifying the results shown in Fig. 1 and highlighting some of the practical merits of DTMs compared to existing approaches. Finally, we conclude by discussing how the capabilities of probabilistic accelerators for machine learning may be scaled by merging them with traditional neural networks (NNs). We demonstrate that such a hybrid model composed of a DTM and a neural network can achieve performance parity with a purely NN-based model on the CIFAR-10 benchmark while using a ten times smaller neural network component (both in terms of the number of parameters and FLOPs) and thus consuming roughly an order of magnitude less energy in total.
@@ I The Challenge with EBMs
Fitting an MEBM corresponds to assigning low energies to values of $x$ where data are abundant and high energies to values of $x$ that are far from data. Real-world data are often clustered into distinct modes [19, 8], meaning that an MEBM that fits data well will have a complex, rugged energy landscape with many deep valleys surrounded by tall mountains. This complexity is illustrated by the cartoon in Fig. 2 (a).
@@
Specifically, sampling algorithms that operate in high dimensions (such as Gibbs sampling [46]) are locally informed iterative procedures, meaning that they sample a landscape by randomly making small movements in the space based on low-dimensional information. When using such a procedure to sample from Eq. (1), the probability that the iteration will move up in energy to some state $X[k+1]$ is exponentially small in the energy increase compared to the current state $X[k]$, i.e.,
@@
For large differences in energy, like those encountered when trying to move between two valleys separated by a significant barrier, this probability can be very close to zero. In standard Markov-chain analyses, such barriers lead to exponentially large expected transition times between modes, which is reflected in a rapidly growing mixing time.
The mixing–expressivity tradeoff (MET) summarizes this issue with existing probabilistic computer architectures, reflecting the fact that modeling performance and sampling hardness are coupled for MEBMs. Specifically, as the expressivity (modeling performance) of an MEBM increases, its mixing time (the amount of computational effort needed to draw independent samples from the MEBM’s distribution) becomes progressively longer, resulting in expensive inference and unstable training [12, 20]; see Appendix G for a definition of mixing time.
@@
Figure 2: The mixing–expressivity tradeoff. (a) A cartoon illustrating the mixing–expressivity tradeoff in EBMs. It shows a projection of an energy landscape fit to a simple dataset. The “airplane” mode is well separated from the “dog” mode, with very little data in between. Progressively better fits of the EBM to the data tend to feature larger energy barriers $\Delta E$ between the modes, making the EBM increasingly difficult to sample from. (b) An example of the effect of the mixing–expressivity tradeoff on model performance as measured using the Fashion-MNIST dataset. The blue curve in the plot shows the results of experiments on MEBMs with limited allowed mixing time. Performance and mixing time are strongly correlated. Mixing times were computed by fitting an exponential function to the large-lag behavior of the autocorrelation function; see Appendix L. In contrast, a DTM (orange cross) has higher performance despite substantially lower sampling requirements.
@@ II Denoising thermodynamic models
DTMs merge EBMs with diffusion models, offering an alternative path for probabilistic computing that mitigates the MET. DTMs are a slight generalization of recent work from deep learning practitioners that has pushed the frontier of EBM performance [26, 71, 70, 72].
@@
In traditional diffusion models, the forward process is made to be sufficiently fine-grained (using a large number of steps $T$) such that the conditional distribution of each step in the reverse process takes some simple form (such as Gaussian or categorical). This simple distribution is parameterized by a neural network, which is then trained to minimize the Kullback–Leibler (KL) divergence between the joint distributions $Q$ and $P_{\theta}$,
@@
See Appendix B.B.2 for more details.
@@
In this parameterization, the dependence on $x^{t}$ enters entirely through the forward energy $\mathcal{E}^{f}_{t-1}(x^{t-1},x^{t})$, which constrains $x^{t-1}$ to stay close to the noisy input $x^{t}$, while $\mathcal{E}^{\theta}_{t-1}$ shapes the local structure of the distribution over $x^{t-1}$.
Equation (7) allows for a compromise between the number of steps in the approximation to the reverse process and the difficulty of sampling at each step. As the number of steps in the forward process is increased, the effect of each noising step becomes smaller, meaning that $\mathcal{E}^{f}_{t-1}$ more tightly binds $x^{t}$ to $x^{t-1}$. This binding can simplify the distribution given in Eq. (7) by imposing an energy penalty that prevents it from being strongly multimodal; see Appendix B.B.4 for further discussion.
@@
Introducing latent variables allows the size and complexity of the probabilistic model to be increased independently of the data dimension. In this generalized form, $\mathcal{E}^{f}_{t-1}$ still enforces proximity between $x^{t-1}$ and $x^{t}$, while $\mathcal{E}^{\theta}_{t-1}$ uses the latent variables $z^{t-1}$ to enrich the local conditional structure without directly depending on $x^{t}$.
@@
See Appendix B.B.6 and Ref. [26] for further details and discussion of this property. Note that it relies on the normalizing constant associated with the distribution in Eq. (6) being independent of $x^{t-1}$.
@@ III Denoising Thermodynamic Computers
Practical implementations of the DTCA utilize natural-to-implement EBMs that exhibit sparse and local connectivity, as is typical in the literature [47]. This constraint allows sampling of the EBM to be performed by massively parallel arrays of primitive circuitry that implement Gibbs sampling. Refer to Appendices C and D for further theoretical discussion of the hardware architecture.
A key feature of the DTCA is that $\mathcal{E}^{f}_{t-1}$ can be implemented efficiently using our constrained EBMs. Specifically, for both continuous and discrete diffusion, $\mathcal{E}^{f}_{t-1}$ can be implemented using a single pairwise interaction between corresponding variables in $x^{t}$ and $x^{t-1}$; see Appendices B.B.1 and D.D.1 for details. This structure can be reflected in how the chip is laid out to implement these interactions without violating locality constraints.
Critically, Eq. (8) places no constraints on the form of $\mathcal{E}_{t-1}^{\theta}$. Therefore, we are free to use EBMs that our hardware implements especially efficiently. At the lowest level, this corresponds to high-dimensional, regularly structured latent-variable EBMs. If more powerful models are desired, these hardware latent-variable EBMs can be arbitrarily scaled by combining them into software-defined graphical models.
@@
Figure 4: A programmable source of random bits. (a) A laboratory measurement of the operating characteristic of our RNG. The probability of the output voltage signal being in the high state ($x=1$) can be programmed by varying an input voltage. The relationship between $\mathbb{P}(x=1)$ and the input voltage is well-approximated by a sigmoid function. The inset shows the output voltage signal as a function of time for different input voltages. (b) The autocorrelation function of the RNG at the unbiased point ($\mathbb{P}(x=1)=0.5$). The decay is approximately exponential with the rate $\tau_{0}\approx 100\text{ns}$. (c) Estimating the effect of manufacturing variation on RNG performance. Each point in the plot represents the results of a simulation of an RNG circuit with transistor parameters sampled according to a procedure defined by the manufacturer’s PDK. Each color represents a different process corner, each for which $\sim 200$ realizations of the RNG were simulated. The “typical” corner represents a balanced case, whereas the other two are asymmetric corners where the two types of transistors (NMOS and PMOS) are skewed in opposite directions. The slow NMOS and fast PMOS case is worst performing for us due to an asymmetry in our design.
@@
Implementing our proposed hardware architecture using Boltzmann machines is particularly simple. A device will consist of nodes connected into a regular grid. Each node represents a single Bernoulli random variable $x_{i}$, implemented as a transistor-based sampling circuit (later referred to as RNG). The bias of a sampling circuit (the probability that it produces 1 as opposed to 1-1) is constrained to be a sigmoidal function of an input voltage, allowing its update probability to be conditioned by the values of its neighbors as given in Eq. (11). This can be implemented using a simple circuit that adds currents such as a resistor network (see Appendix E.E.1).
@@
Specifically, the EBMs employed in this work were sparse, deep Boltzmann machines comprising $L\times L$ grids of binary variables, where $L=70$ was used in most cases. Each variable was connected to several (in most cases, 12) of its neighbors following a simple pattern. At random, some of the variables were selected to represent the data $x_{t-1}$, and the rest were assigned to the latent variables $z_{t-1}$. Then, an extra node was connected to each data node to implement the coupling to $x_{t}$. See Appendix D for further details on the Boltzmann machine architecture.
Due to our chosen connectivity patterns, our Boltzmann machines are bipartite (two-colorable), meaning that nodes can be separated into two blocks, such that no two nodes within the same block are connected to each other. Since each color block can be sampled in parallel, a single iteration of Gibbs sampling corresponds to first sampling one color block conditioned on the other and then swapping their roles. Therefore, the total duration of one full Gibbs iteration (so updating all nodes) takes roughly $2\tau_{0}$ in wall-clock time, where $\tau_{0}$ is the decorrelation time of an individual RNG, that is the time it takes for a sampling circuit inside a single node of the Boltzmann machine to equilibrate. It should be noted that $\tau_{0}$ is set by the intrinsic stochastic dynamics of the subthreshold circuit and its operating point. It is not a freely tunable algorithmic parameter, but can only be modified indirectly through circuit redesign.
Starting from some random initialization, this block-sampling procedure can be repeated for $K$ iterations (where $K$ is longer than the mixing time of the sampler, typically $K\approx 1000$) to draw samples from Eq. (7) for each step in the approximation to the reverse process. For a DTM running on a DTCA chip, the total time required to draw a single sample is therefore proportional to $TK\,\tau_{0}$, where $T$ is the number of denoising steps (i.e., EBMs that the DTM is comprised of). It is important to distinguish $\tau_{0}$ from $K_{\text{mix}}$. The former is an entirely hardware-related parameter, while the latter is the number of Gibbs iterations that the abstract Boltzmann machine (irrespective of hardware implementation) needs to approximately converge to its equilibrium distribution.
@@
To enable a near-term, large-scale realization of the DTCA, we leveraged the shot-noise dynamics of subthreshold transistors [25] to build an RNG that is fast, energy-efficient, and small. Our all-transistor RNG is programmable and has the desired sigmoidal response to a control voltage, as shown by experimental measurements in Fig. 4 (a). The stochastic voltage signal output from the RNG has an approximately exponential autocorrelation function that decays in approximately $100$ ns, as illustrated in Fig. 4 (b). As shown in Ref. [25], this time constant is much larger than the lower limit imposed by the correlation time of the noise in our transistors. The RNG could, therefore, be made much faster via an improved design. Appendix K provides further details about our RNG.
@@
In practice, semiconductor manufacturing is never perfectly precise. Each fabrication step introduces small fluctuations in transistor characteristics, producing global shifts across wafers as well as random differences between neighboring devices on the same chip. Because these variations can change how fast or efficiently a circuit operates, it is important to verify that our approach remains reliable under realistic fabrication conditions. A practical advantage of our all-transistor RNG is that detailed and proven foundry-provided models can be used to study the effect of manufacturing variations on our circuit design. In Fig. 4 (c), we use this process development kit (PDK) to study the speed and energy consumption of our RNG as a function of both systematic inter-wafer skews to the transistor parameters (process corners) and the expected variation within a single chip. We find that the RNG works reliably despite these non-idealities, meaning it can be readily scaled to the massive grids required by the DTCA.
@@
where $E_{\text{rng}}$ comes from the data in Fig. 4 (c). The term $E_{\text{bias}}$ is estimated using a physical model of a possible biasing circuit, and $E_{\text{clock}}$ and $E_{\text{comm}}$ are derived from physical reasoning about the costs of the clock and inter-cell communications, respectively. $K_{\text{mix}}$ is the number of sampling iterations required to satisfactorily mix the chain for inference, which is generally less than the number of iterations used during training. $K_{\text{mix}}=250$ was used for the DTM (see Appendix E.E.4), while the mixing time measured in Fig. 2 was used for the MEBM.
This model is approximate, but it captures the underlying physics of a real device and provides a reasonable order-of-magnitude estimate of the actual energy consumption. Generally, given the same transistor process we used for our RNG and some reasonable selections for other free parameters of the model, we can estimate $E_{\text{cell}}\approx 2\;\text{fJ}$. See Appendix E for an exhaustive derivation of this model.
We use a simple model for the energy consumption of the GPU that underestimates the actual values. We compute the total number of floating-point operations (FLOPs) required to generate a sample from the trained model and divide that by the FLOP/joule specification given by the manufacturer. See Appendix F for further discussion.
Figure 5: Detailed results on the Fashion-MNIST dataset. (a) Images generated by a denoising model. Here, to achieve better-looking images, several binary variables were combined to represent a single grayscale pixel. The noisiness of the grayscale levels is an artifact of our embedding method; see Appendix I. (b) An experiment showing how DTMs are more stable to train than MEBMs. Complementing DTMs with the ACP completely stabilizes training. For the DTMs, the maximum $r_{yy}[K]$ value over all the layers is shown. (c) The effect of scaling EBM complexity on DTM performance. The grid size $L$ was modified to change the number of latent variables compared to the (fixed) number of data variables. Generally, EBM layers with more connectivity and longer allowed mixing times can utilize more latent variables and, therefore, achieve higher performance.